OpenStudy (anonymous):

Find the slope and y-intercept: -11-8y=0

OpenStudy (freckles):

is that really the equation?

OpenStudy (marcelie):

arent you missing an x ? in order for the equation look like y=mx+b

OpenStudy (anonymous):

Yes

OpenStudy (freckles):

-11-8y=0 add 8y on both sides -11=8y divide both sides by 8 -11/8=y this tells you the y-intercept and since it is a horizontal line the slope is...

OpenStudy (anonymous):

M=0. Thank you!

OpenStudy (freckles):

yes the slope is 0 and the y-intercept is...

OpenStudy (freckles):

y=a says all the points that are on the horizontal line y=a has y coordinate a so that means the y-intercept has to be a.